# Fast Transaction Scheduling in Blockchain Sharding

### v1

@@ 4.1. Analysis for General Graph
For the general graph, the processing time of any schedule for ${\mathcal{T}}$ in a graph $G_{\mathcal{T}}$ is at least $\Omega(l)$, where $l$ is the maximum number of transactions that access any object $o_{j}\in{\mathcal{O}}$.
Proof.
@@
Let ${\mathcal{T}}_{o_{j}}$ denote the set of transactions that use object $o_{j}\in{\mathcal{O}}$. Let $l_{o_{j}}=|{\mathcal{T}}_{o_{j}}|$, and $l=\max_{j}l_{o_{j}}$ be the maximum degree in the graph for any object in ${\mathcal{O}}$. Then, the length of the schedule is at least $l$ because an object has to be accessed once at a time by $l$ transactions. Consequently, the processing time of any schedule for ${\mathcal{T}}$ is at least $\Omega(l)$. ∎
Lemma 0 (Upper Bound).
@@
Proof.
The maximum degree in the conflict graph $G_{\mathcal{T}}$ is at most $kl$, since each transaction accesses $k$ objects and each object is accessed by at most $l$ transactions. Therefore, the conflict graph $G_{\mathcal{T}}$ can be greedily colored with at most $\xi\leq kl+1$ colors.
@@
Theorem 3.
For the general graph, Algorithm 1 gives an $O(kd)$ approximation to the optimal schedule.
Proof.
From Lemma 2, the time to process all the transactions is at most $O(kld)$. Moreover, from Lemma 1, $\Omega(l)$ is the lower bound, which gives us $O(kd)$ approximation to the optimal schedule.
4.2. Analysis for Specialized Graphs
@@
Let $G_{s}$ be a hypercube graph (Leighton, 2014). Then, there is a connecting path between each pair of shards with length $\log s$; hence $D=\log s$, and $d\leq D=\log s$. Therefore, Theorem 3 implies that there is an execution schedule which is a $O(k\log s)$ approximation of the optimal schedule.
Similar to the Hypergraph (Leighton, 2014), we can use the same result in the Butterfly networks (Leighton, 2014) and $g$-dimensional grids (Chan, 1989), where $g=O(\log s)$, which also gives an approximation for execution time as $O(k\log s)$.
@@
Corollary 0.
@@
5. Centralized Scheduler with Buckets
In the Centralized scheduler (Algorithm 1), a single shard (the leader shard) colors the transactions and determines the schedule. We propose an alternative algorithm that uses buckets to organize transactions based on the distances between the transaction and its accessed objects. This approach offers improved bounds for special cases when the $k$ objects are selected randomly, as well as for the Line graph.
Here, we distribute the coloring load in different shards, i.e., each bucket has its own leader shard, which is known to all the shards and determines the schedule to process the transaction of that bucket. We divide the transactions into buckets $B_{0},B_{1},\ldots$, according to the maximum distance between the home shard of the transaction and its accessing objects (i.e., destination shards). In bucket $B_{i}$, we add transactions whose maximum distance between home shard and its destination shard is in range $[2^{i},2^{i+1})$. We compute a schedule for each bucket. We also prioritize the processing of the lower index bucket transactions over the higher index bucket so that we can get the locality benefits and reduce transaction latency, that is, first, we calculate a schedule and process transactions of $B_{0}$, then for $B_{1}$, then for $B_{2}$, and so on.
The pseudocode of the centralized scheduler with the bucketing approach is presented in Algorithm 2. We assume that shard graph $G_{s}$, a set of transactions ${\mathcal{T}}$ that need to be processed, are given as input to the algorithm. Then, the scheduler provides the schedule and processes all the transactions ${\mathcal{T}}$. We assume that each bucket $B_{i}$ has its own leader shard $\ell_{i}$, which is known to all shards in $G_{s}$. First, each transaction determines its bucket, and then we run Algorithm 1 for each bucket by providing the shard graph $G_{s}$, set of transactions of each bucket $B_{i}$ and leader of that bucket as parameters.
@@ 5. Centralized Scheduler with Buckets
4 Each Bucket $B_{i}$ has its own leader shard $\ell_{i}$ which is known to all shards in $G_{s}$;
@@ 8       Run CentralizedScheduler($G_{s}$, $B_{i}$, $\ell_{i}$);
5.1. Analysis for General Graph with Objects Chosen Randomly
Suppose each shard contains one transaction that needs to be processed and each transaction accesses at most $k$ objects picked randomly and uniformly from different shards. We use the following version of Chernoff bound in our analysis.
Lemma 0 (Chernoff bound).
@@ 5.1. Analysis for General Graph with Objects Chosen Randomly
Proof.
Consider an object $o_{z}\in{\mathcal{O}}$. Let $X^{(z)}_{i}$ be the random variable such that $X^{(z)}_{1}=1$ if transaction $T_{i}$ accesses object $o_{z}$. Let $Y_{z}=\sum_{j=i}^{s}X^{(z)}_{j}$, and let $\mu=E[Y_{z}]$ be the number of transactions that access object $o_{z}$.
@@
From Lemma 1 (Chernoff bound) for a random variable $Y_{z}$ and any $\delta>1$ with expected value $\mu$ is:
@@
In our case, $Y_{z}$ represents the number of transactions accessing a specific object, and $\mu=k$ (the expected number of transactions accessing each object).
We want to find the probability that the number of transactions accessing this object is greater or equal to $c$, i.e., $\mathrm{Pr}(Y_{z}\geq c)$. To apply the Chernoff bound, we set $c=(1+\delta)k$, which gives us $\delta=\frac{c-k}{k}$ and $\mu=k$. Let’s substitute these into the Chernoff bound:
@@
We want this probability to be less than or equal to $\frac{1}{s}$, so:
$s\cdot e^{-\frac{(c-k)}{2}}\leq\frac{1}{s}\ .$
Solving for $c$, we get:
$c\geq k+4\ln s\ .$
We assume $c=(1+\delta)k$, where $\delta>1$. Thus, we can take
$c\geq 2k+4\ln s\ .$
Therefore, $l\leq 2k+4\ln s$ with probability at least $1-\frac{1}{s}$.
Lemma 0 (Lower Bound).
@@
Lemma 0 (Upper Bound).
@@
If each transaction accesses at most $k$ objects randomly from different shards, then the time to process the transactions in any bucket $B_{i}$ is at most $O(2^{i}k\log D\cdot(k+\log s))$.
Proof.
@@
Suppose $|B_{i}|$ is the total number of transactions in the bucket $B_{i}$. Let ${\mathcal{T}}_{x}^{b_{i}}$ denote the set of transactions in $B_{i}$ that use object $o_{j}\in{\mathcal{O}}$. Let $l_{x}=|{\mathcal{T}}_{x}^{b_{i}}|$ and $l_{i}=max_{x}l_{x}$ be the maximum number of transactions that used any object in ${\mathcal{O}}$. From Lemma 2, if the transaction accesses the objects randomly, the maximum degree of any object is bounded by $2k+4\ln s$. Thus, we can write $l_{i}\leq 2k+4\ln s$. Since all the transactions in the bucket $B_{i}$ access at most $k$ objects and each object is used by at most $l_{i}$ transactions. Thus, the maximum degree in the transaction graph of bucket $B_{i}$ is at most $kl_{i}$, and therefore, the conflict graph $G_{{\mathcal{T}}}$ can be greedily colored with $\xi\leq kl_{i}+1$ colors.
@@
The distance between a transaction and its accessing object of bucket $B_{i}$ is at most $2^{i+1}-1<2^{i+1}$. Moreover, the simulation of Algorithm 1 has at most $4$ rounds of interactions between the transaction and its accessing objects for the confirmation and commit. Therefore, each color corresponds to $4\cdot 2^{i+1}$ rounds. Thus, it takes $4\cdot 2^{i+1}(kl_{i}+1)=O(2^{i}kl_{i})=O(2^{i}k(k+\log s))$ rounds to confirm and commit transactions.
Thus, if we have only one bucket $B_{i}$, then the time to process all the transactions of the bucket $B_{i}$ is at most:
(1) Maximum processing time for bucket $\text{Maximum processing time for bucket $B_{i}$}=2^{i+3}(kl_{i}+1)$
In our algorithm, transactions from lower index buckets are prioritized for processing over higher index buckets. As we have at most $\lfloor\log D\rfloor+1$ buckets, we analyzed the effect of it to process all the transactions.
Assume bucket $B_{i}$ is the worst bucket, and it takes more time to process all the transactions on its bucket than any other bucket. From equation 1, if we only have bucket $B_{i}$, the cost required for the scheduler to process all the transactions is at most $B_{i}^{c}=4\cdot 2^{i+1}(kl_{i}+1)$ rounds. However, if we consider all the buckets, the scheduler needs to process transactions from bucket 00 to bucket $\log D$. Thus, the total number of rounds needed to process all bucket transactions is at most:
(2) $\sum_{x=o}^{\log D}B_{x}^{c}\leq 2B_{i}^{c}\log D=2\cdot 4\cdot 2^{i+1}(kl_{i}+1)\log D=2^{i+4}(kl_{i}+1)\log D$
Thus, the total processing time to process all the transactions from all the buckets is at most $2^{i+4}(kl_{i}+1)\log D=O(2^{i}kl_{i}\log D)=O(2^{i}k\log D\cdot(k+\log s))$. ∎
From Equation 1, if we consider only one bucket $B_{i}$ and $B_{i}$ is the worst bucket, then total transactions processing time is at most $2^{i+3}(kl_{i}+1)$. However, from Equation 2, if we consider all the buckets, then total transactions processing time is at most $2^{i+4}(kl_{i}+1)\log D$. Thus, if we use a bucketing approach, then we need to pay an extra $O(\log D)$ cost over the worst bucket. Hence, we have the following corollary.
Corollary 0.
If we use a bucketing approach, we need to pay an extra $O(\log D)$ cost over the worst bucket.
Theorem 6.
@@
For the General graph, if the transaction accesses $k$ objects randomly from different shards, then Algorithm 2 gives an $O(k\log D\cdot(k+\log s))$ approximation to the optimal schedule with probability at least $1-1/s$.
Proof.
From Lemma 4, the time to process all the transactions in the bucket $B_{i}$ is at most $O(2^{i}k\log D\cdot(k+\log s))$. Moreover, from Lemma 3, $2^{i}$ is the lower bound, which gives us $O(k\log D\cdot(k+\log s))$ approximation to the optimal schedule. ∎
5.2. Analysis for Line Graph with Buckets
We consider a graph as a line graph with $(V,E)$ where shards are connected in a line. We consider the sequence of $s$ shards, where $V=\{v_{1},v_{2},v_{3},\cdots,v_{s}\}$, and there is a edge between two vertices $v_{i},v_{i+1}$ of weight $1$ for $11\leq i\leq s-$ 1. We consider the position of the shards in the line graph to be left to right. Thus, $v_{1}$ is the leftmost shard, and $v_{s}$ is the rightmost shard. We can also consider mapping the line graph into the clique graph we discussed in the previous section. That means we can map every shard of the line to the corresponding shard in a clique graph, and every edge of a line graph corresponds to some edge in a clique graph, which connects direct neighbors in the line graph. Additional edges in the clique graph that connect indirect neighbors in the line graph have weights equal to the distance between the neighbors.
Lemma 0 (Lower Bound).
For the Line graph, the time to process the transactions in bucket $B_{i}$ is at least $2^{i}+\frac{l_{i}^{2}}{4}$, where $l_{i}$ is the maximum number of transactions in $B_{i}$ that access any object $o_{j}\in{\mathcal{O}}$.
Proof.
Suppose shards are arranged in a line, and each shard contains one transaction to be processed. Let $|B_{i}|$ be the total number of transactions in the bucket $B_{i}$. Suppose there is an object $o_{j}\in{\mathcal{O}}$, which is accessed by all the transactions that are in $B_{i}^{\prime}\subseteq B_{i}$. Thus the maximum number of transactions $l_{i}$ in $B_{i}$ that access an object $o_{j}$ is $|B_{i}^{\prime}|$, i.e. $l_{i}=|B_{i}^{\prime}|$.
We can decompose the bucket $B_{i}^{\prime}$ into partial buckets $\bar{B}^{\prime}_{0,i},\bar{B}^{\prime}_{1,i},\cdots,\bar{B}^{\prime}_{z,i}$ where partial bucket $\bar{B}^{\prime}_{z,i}$ store the transaction which access object $o_{j}$ from distance $[2^{z},2^{z+1})$.
Without loss of generality, suppose the object $o_{j}$ is in the middle of the line. Since each shard hosts a single transaction, the size of the bucket is bounded by the distance that comprised transactions have to traverse $|B_{z,i}^{\prime}|<2^{i+1}$. Thus $z\leq\log l_{i}$. Now calculate the number of transactions that are at a distance of at least $z-2\Longrightarrow\log l_{i}-2$.
$\sum_{z=0}^{\log l_{i}-2}|B_{z,i}^{\prime}|<\sum_{z=0}^{\log l_{i}-2}2^{z}\leq 2\cdot 2^{\log l_{i}-2}\leq l_{i}/2\ .$
As the transactions are in line, the remaining $l_{i}/2$ transactions must be in the bucket index $\bar{B}^{\prime}_{>z-2,i}$. Consider a bucket $\bar{B}^{\prime}_{z-1,i}$ which stores the transaction accessing object $o_{j}$ from distance $22^{z-1}\Longrightarrow 2^{\log l_{i}-1}=l_{i}/$ 2. This implies that other $l_{i}/2$ transactions are at least $l_{i}/2$ distance from object $o_{j}$. Thus, the maximum degree for the object $o_{z}$ is at least $l_{i}/2\cdot l_{i}/2=\frac{l_{i}^{2}}{4}$ because each $l_{i}/2$ conflicting transaction needs to be serialized, and they are far from object $o_{j}$ and need to be travel at least $l_{i}/2$ distance.
Moreover, there is at least one transaction in bucket $B_{i}$, which accesses the object at a distance of at least $2^{i}$; that is why they are in bucket $B_{i}$. Consequently, the total time to process all the transactions of bucket $B_{i}$ is at least $2^{i}+\frac{l_{i}^{2}}{4}$. ∎
Lemma 0 (Upper Bound).
For the Line graph, the time to process the transactions in bucket $B_{i}$ is at most $O(2^{i}kl_{i}\log D)$, where $l_{i}$ is the maximum number of transactions in $B_{i}$ that access any object $o_{j}\in{\mathcal{O}}$.
Proof.
Suppose shards are arranged in a line and consist of transactions that need to be processed in each shard. Let $|B_{i}|$ be the total number of transactions in the bucket $B_{i}$. In the worst case, all of the transactions can access the same object $o_{j}$. Thus, the maximum degree of the object $o_{j}$ access by transactions in bucket $B_{i}$ is $l_{i}=|B_{i}|$. Since all the transactions in the bucket $B_{i}$ access at most $k$ objects and each object is used by at most $l_{i}$ transactions. The maximum degree of the transaction graph of bucket $B_{i}$ is at most $kl_{i}$, and therefore, the transaction graph can be greedily colored with $\xi\leq kl_{i}+1$ colors.
The distance between a transaction and its accessing objects of bucket $B_{i}$ is at most $2^{i+1}-1<2^{i+1}$. Moreover, for each color, Algorithm 1 takes $4\cdot 2^{i+1}$ rounds to confirm and commit the transactions. Thus, transactions of the bucket $B_{i}$ take at most $4\cdot 2^{i+1}(kl_{i}+1)=O(2^{i}kl_{i})$ rounds to process transactions.
@@ 5.2. Analysis for Line Graph with Buckets
Additionally, from Corollary 5, if we consider all the buckets, we need to pay an extra $O(\log D)$ cost. Thus, the total execution time to process all the transactions of the bucket $B_{i}$ is at most $O(2^{i}kl_{i}\log D)$. ∎
Theorem 9 (Approximation).

### v3

@@ IV-A Analysis for General Graph
For graph $G_{s}$, the processing time of any schedule for ${\mathcal{T}}$ in a graph $G_{\mathcal{T}}$ is at least $l$, where $l$ is the maximum number of transactions that access any object $o_{j}\in{\mathcal{O}}$.
Proof:
@@
Let ${\mathcal{T}}_{o_{j}}$ denote the set of transactions that use object $o_{j}\in{\mathcal{O}}$. Let $l_{o_{j}}=|{\mathcal{T}}_{o_{j}}|$, and $l=\max_{j}l_{o_{j}}$ be the maximum number of transactions that access any object in ${\mathcal{O}}$. Then, the length of the schedule is at least $l$ because an object has to be accessed once at a time by $l$ transactions. Consequently, the processing time of any schedule for ${\mathcal{T}}$ is at least $l$. ∎
Lemma 2 (Upper Bound)
@@
Proof:
The maximum degree in the conflict graph $G_{\mathcal{T}}$ is at most $kl$, since each transaction accesses at most $k$ objects and each object is accessed by at most $l$ transactions. Therefore, the conflict graph $G_{\mathcal{T}}$ can be greedily colored with at most $\xi\leq kl+1$ colors.
@@
Theorem 1
For the general graph $G_{s}$, Algorithm 1 gives an $O(kd)$ approximation to the optimal schedule.
Proof:
From Lemma 2, the time to process all the transactions is at most $O(kld)$. Moreover, from Lemma 1, $l$ is the lower bound, which gives us $O(kd)$ approximation to the optimal schedule.
@@
Let $G_{s}$ be a Hypercube graph [48]. Then, there is a connecting path between each pair of shards with length $\log s$; hence $D=O(\log s)$, and $d\leq D=O(\log s)$. Therefore, Theorem 1 implies that there is an execution schedule which is a $O(k\log s)$ approximation of the optimal schedule. A similar property holds for the Butterfly, and $g$-dimensional grids graph [49], where $g=O(\log s)$. Thus, we have:
Corollary 2
@@
V Centralized Scheduler with Buckets
In the Centralized scheduler (Algorithm 1), a single shard (the leader shard) colors the transactions and determines the schedule. We propose an alternative algorithm that uses buckets to organize transactions based on the distances between the transaction and its accessed objects. This approach offers improved bounds for special cases when the transaction accesses $k$ objects randomly, as well as for the Line graph.
Here, we distribute the coloring load in different shards, i.e., each bucket has its own leader shard, which is known to all the shards and determines the schedule to process the transactions of that bucket. We divide the transactions into buckets $B_{0},B_{1},\ldots$, according to the maximum distance between the home shard of the transaction and its accessing objects (i.e., destination shards). In bucket $B_{i}$, we add transactions whose maximum distance between home shard and its destination shard is in range $[2^{i},2^{i+1})$. We compute a schedule for each bucket. We also prioritize the processing of the lower index bucket transactions over the higher index bucket so that we can get the locality benefits and reduce transaction latency. That is, first, we calculate a schedule and process transactions of $B_{0}$, then for $B_{1}$, then for $B_{2}$, and so on.
The pseudocode of the centralized scheduler with the bucketing approach is presented in Algorithm 2. We assume that shard graph $G_{s}$, a set of transactions ${\mathcal{T}}$ that need to be processed, are given as input to the algorithm. Then, the scheduler provides the schedule and processes all the transactions ${\mathcal{T}}$. We assume that each bucket $B_{i}$ has its own leader shard $S_{\ell_{i}}$, which is the home shard of some transaction in $B_{i}$ and is known to all shards in $G_{s}$. First, each transaction determines its bucket, and then we run Algorithm 1 for each bucket by providing the shard graph $G_{s}$, set of transactions of each bucket $B_{i}$ and leader of that bucket as parameters. After completing the transaction processing of bucket $B_{i}$, the leader of $B_{i}$ informs the leader of $B_{i+1}$ to run the Algorithm 1 and process their transactions.
@@ V Centralized Scheduler with Buckets
4 Each Bucket $B_{i}$ has its own leader shard $S_{\ell_{i}}$ which is the home shard of some transaction in $B_{i}$, and it is known to all shards in $G_{s}$;
@@ 8       Run CentralizedScheduler($G_{s}$, $B_{i}$, $S_{\ell_{i}}$);
V-A Analysis for General Graph with Random Object Accesses
Suppose each shard contains one transaction that needs to be processed, and each transaction accesses at most $k$ objects picked randomly from different shards. We consider the case where there is one object per shard, and each shard is the home of one transaction, that is, there are $s$ objects and transactions. We use the following version of Chernoff bound in our analysis.
Lemma 3 (Chernoff bound)
@@ V-A Analysis for General Graph with Random Object Accesses
Proof:
Consider an object $o_{z}\in{\mathcal{O}}$. Let $X^{(z)}_{i}$ be the random variable such that $X^{(z)}_{i}=1$ if transaction $T_{i}$ accesses object $o_{z}$. Let $Y_{z}=\sum_{i=1}^{s}X^{(z)}_{i}$, and let $\mu=E[Y_{z}]$ be the expected number of transactions that access object $o_{z}$.
@@
From Lemma 3 (Chernoff bound) for a random variable $Y_{z}$ and any $\delta>1$ with expected value $\mu=k$ we have:
@@
We want to find the probability that the number of transactions accessing this object is at least $c$, i.e., $\mathrm{Pr}(Y_{z}\geq c)$. To apply the Chernoff bound, we set $c=(1+\delta)k$, which gives us $\delta=\frac{c-k}{k}$. We substitute these into the Chernoff bound:
@@
We want this probability to be less than or equal to $\frac{1}{s}$, so $s\cdot e^{-\frac{(c-k)}{2}}\leq\frac{1}{s}.$. Solving for $c$, we get: $c\geq k+4\ln s$. We assume $c=(1+\delta)k$, where $\delta>1$. Thus, we can take $c\geq 2k+4\ln s$. Therefore, $l\leq 2k+4\ln s$ with probability at least $1-\frac{1}{s}$.
Lemma 5 (Lower Bound)
@@
Suppose $|B_{i}|$ is the total number of transactions in the bucket $B_{i}$. Let $B_{i}(o_{x})\subseteq B_{i}$ denote the set of transactions in $B_{i}$ that use object $o_{x}\in{\mathcal{O}}$. Let $l_{i}=\max_{o_{x}\in{\mathcal{O}}}|B_{i}(o_{x})|$ be the maximum number of transactions in $B_{i}$ that used any object in ${\mathcal{O}}$.
Lemma 6 (Upper Bound)
@@
If each transaction accesses at most $k$ objects randomly from different shards, then the time to process the transactions in any bucket $B_{i}$ is at most $2^{i+3}(k(2k+4\ln s)+1)$, with probability at least $1-1/s$.
Proof:
From Lemma 4, if a transaction picks to access $k$ objects randomly, then an object is used by at most $2k+4\ln s$ transactions (with probability at least $1-1/s$). Thus, we can write $l_{i}\leq 2k+4\ln s$.
@@
Let $G_{B_{i}}$ denote the transaction conflict graph for the transactions in bucket $B_{i}$. Since all the transactions in the bucket $B_{i}$ access at most $k$ objects and each object is used by at most $l_{i}$ transactions, the maximum degree in the transaction graph of bucket $B_{i}$ is at most $kl_{i}$. Therefore, the conflict graph $G_{B_{i}}$ can be greedily colored with $\xi\leq kl_{i}+1$ colors.
@@
For a transaction in $B_{i}$, the distance to its accessing object is at most $2^{i+1}-1<2^{i+1}$. Moreover, the execution of Algorithm 1 has at most $4$ rounds of interactions between the transaction and its accessing objects for the confirmation and commit. Therefore, each color corresponds to at most $4\cdot 2^{i+1}$ rounds. Thus, it takes at most
$4\cdot 2^{i+1}\xi\leq 4\cdot 2^{i+1}(kl_{i}+1)\leq 2^{i+3}(k(2k+4\ln s)+1)$
rounds to confirm and commit transactions. ∎
Theorem 2
@@
For the General graph, if each transaction accesses $k$ objects randomly from different shards, then Algorithm 2 gives an $O(k\log D\cdot(k+\log s))$ approximation to the optimal schedule with probability at least $1-1/s$.
Proof:
@@
Let $B_{m}$ be the bucket that takes the most time to process all its transactions. Thus, when we consider all the $\lfloor\log D\rfloor+1$ buckets, from Lemma 6, the total number of rounds needed to process all transactions is at most:
$\sum_{i=0}^{\lfloor\log D\rfloor+1}2^{i+3}(k(2k+4\ln s)+1)$
$\leq 2^{m+3}(k(2k+4\ln s)+1)(\log D+2)\ .$
From Lemma 5, $2^{m}$ is a lower bound on the cost of processing the transactions in bucket $B_{m}$. Therefore, the algorithm approximates the optimal cost within a factor
$\frac{2^{m+3}(k(2k+4\ln s)+1)(\log D+2)}{2^{m}}=O((k^{2}+\log s)\log D)\ .$
V-B Analysis for Line Graph with Buckets
We consider graph $G_{s}=(V,E)$ which is a line graph where shards are connected in the sequence of $s$ shards, $S_{1},S_{2},\cdots S_{s}\in V$, where any two consecutive shards $S_{i},S_{i+1}$ are connected with an edge of weight $1$, for $11\leq i\leq s-$ 1. Generally, for any pair $S_{i},S_{j}$, where $i<j$, there is an edge of weight $j-i$, where $1\leq i,j\leq s$. We consider the orientation of the shards in the line to be $S_{1}$ the leftmost and $S_{s}$ the rightmost. We continue with the analysis where we use the buckets as defined earlier, and we assume that each shard has one transaction to be processed.
Lemma 7 (Lower Bound)
For the Line graph, the time to process the transactions in a non-empty bucket $B_{i}$ is at least $2^{i}+\frac{l_{i}^{2}}{8}$.
Proof:
Let $o_{j}$ be the object for which the parameter $l_{i}$ is maximized. Let $B_{i}^{\prime}=B_{i}(o_{j})$, with $l_{i}=|B_{i}^{\prime}|$. We can decompose the bucket $B_{i}^{\prime}$ into partial buckets $\bar{B}^{\prime}_{0,i},\bar{B}^{\prime}_{1,i},\cdots,\bar{B}^{\prime}_{z,i}$ where partial bucket $\bar{B}^{\prime}_{z,i}$ contains the transactions which access object $o_{j}$ from distance $[2^{z},2^{z+1})$.
Without loss of generality, suppose the object $o_{j}$ is in the middle of the line (assuming $s$ is odd). Since each shard hosts a single transaction, we have $|\bar{B}_{z,i}^{\prime}|\leq 2^{z+1}-2^{z}=2^{z}$. The number of transactions up to $z^{*}=\lceil(\log l_{i})-3\rceil$ (logarithm is base 2) is
$\sum_{z=0}^{\lceil(\log l_{i})-3\rceil}|\bar{B}_{z,i}^{\prime}|\leq\sum_{z=0}^{(\log l_{i})-2}2^{z}\leq 2\cdot 2^{(\log l_{i})-2}\leq\frac{l_{i}}{2}\ .$
As the transactions are in line, the remaining $l_{i}/2$ transactions must be in the bucket index $\bar{B}^{\prime}_{>z^{*},i}$. Bucket $\bar{B}^{\prime}_{z^{*}+1,i}$ has transactions at distance at least $42^{z^{*}+1}\geq 2^{\lceil(\log l_{i})-3\rceil+1}\geq 2^{(\log l_{i})-2}\geq l_{i}/$ 4. This implies that these $l_{i}/2$ transactions are at least $l_{i}/4$ distance from object $o_{j}$.
Thus, $l_{i}/2\cdot l_{i}/4=\frac{l_{i}^{2}}{8}$ is a lower bound on the execution time because each of these $l_{i}/2$ transactions need to be serialized (because the conflict on each other on $o_{j}$), and they also need to travel at least $l_{i}/4$ distance to reach object $o_{j}$.
Moreover, there is at least one transaction in bucket $B_{i}$, which accesses the object at a distance of at least $2^{i}$; that is why they are in bucket $B_{i}$. Consequently, the total time to process all the transactions of bucket $B_{i}$ is at least $2^{i}+\frac{l_{i}^{2}}{8}$. ∎
The distance between a transaction in $B_{i}$ and its accessing objects is at most $2^{i+1}-1<2^{i+1}$. Thus, from Lemma 2, since $d\leq 2^{i+1}$, we get the following result.
Lemma 8 (Upper Bound)
For the Line graph, the time to process the transactions in bucket $B_{i}$ is at most $O(2^{i}kl_{i})$.
Theorem 3 (Approximation)